use Lagrange’s Multiplier method to find the maximum and minimum of f(x,y)= y^2 - x^2 subjected to the constraint of 1/4 x^2 + y^2 = 1


1
Expert's answer
2021-06-25T03:32:25-0400

Given 

"f(x,y) =y^2-x^2"


Let  

"g(x,y) =\dfrac{1}{4} x^{2}+y^{2}-1"

 and let 

"F = f(x,y)+\lambda g(x,y)"


where "\lambda"  is the Lagrangian multiplier.


"F = y^2-x^2 + \lambda(\dfrac{1}{4} x^{2}+y^{2}-1)"


To find the maximum and minimum values, we have to solve the system


"\dfrac{\partial F}{\partial x} = 0"




"\dfrac{\partial F}{\partial y} = 0"

 


"\dfrac{\partial F}{\partial \lambda} = 0"


Hence


"-2x + \dfrac{1}{2}x\lambda = 0"




"2y+2y\lambda=0"

"\dfrac{1}{4} x^{2}+y^{2}-1=0"

If "x=0," we get


"x=0"

"\lambda=-1"


"y^2=1"

When "\lambda=-1,"


"x=0, y=-1" or "x=0, y=1."


If "y=0," we get


"\lambda=4"

"y=0"


"x^2=4"

When "\lambda=4,"


"x=-2, y=0" or "x=2,y=0."


If "x\not=0, y\not=0," we get


"\lambda = 4"




"\lambda=-1"

"\dfrac{1}{4} x^{2}+y^{2}-1=0"

The system has no solution.



"f(0, -1)=(-1)^2-(0)^2=1"

"f(0, 1)=(1)^2-(0)^2=1"


"f(-2, 0)=(0)^2-(-2)^2=-4"

"f(2, 0)=(0)^2-(2)^2=-4"

The maximum and minimum values of f(x,y) are respectively,

maximum value:


"f(0, -1)=f(0, 1)=1"

minimum value:


"f(-2, 0)=f(2, 0)=-4"